### v2

@@ 6 Conclusion
The debonding of a rock during dissolution can lead the material to failure as mechanical properties weaken and the state of stress evolves. This effect can have important implications for fault nucleation and reactivation for underground storage projects. The influence of several parameters such as the degree of cementation, the confining pressure, the initial value of $k_{0}$ and the history of load have been investigated in this paper.
It has been shown that an attractor configuration exists and the grain reorganization occurs during the debonding to reach this state. Especially, the $k_{0}$ evolves to its attractor value, increasing or decreasing. It appears that a $k_{0}$ reduction, which is the most likely in a reservoir configuration, can be at the origin of fault reactivation and therefore induced seismicity.
Two main mechanisms have been emphasized for the grain reorganization, depending on the initial state of stress. The chain forces need to be sorted into two kinds: the unstable and the stable chain forces. The unstable ones support the force thanks to the cementation, when the debonding occurs they collapse. On the opposite, the stable ones stay in place, even if they are no cementation. It appears that the unstable ones are dominant in the case $k_{0}^{init}\leq k_{0}^{attr}$ and the stable ones are dominant in the case $k_{0}^{init}\geq k_{0}^{attr}$. In the latter, the grain reorganization occurs with the grain softening.
7 Software
The Discrete Element Model is solved using the YADE open source software [44]. Some examples of scripts used are available on GitHub at the following link:
8 Acknowledgements
@@ Appendix A The lognormal distribution
It appears this distribution reproduces accurently experimental observations [24]. Indeed, smaller bond surfaces and larger ones can be considered, both are important to the mechanical behavior of the sample. Examples of lognormal distribution used in this paper are given in Figure 19.

### v3

@@ 6 Conclusion
The debonding of a rock during weathering can result in the material failure as a consequence of the deterioration of its mechanical properties, and the subsequent evolution of the state of stress. This effect has the potential to have significant implications for fault nucleation and reactivation in the context of underground storage projects. This paper explores the impact of several parameters such as the degree of cementation, the confining pressure, the initial value of $k_{0}$, and the loading history.
It has been demonstrated that an attractor configuration exists and that grain reorganization occurs during the debonding in order to reach this state. In particular, $k_{0}$ evolves to its attractor value, exhibiting either an increase or a decrease. It appears that a $k_{0}$ reduction, which is the most likely in a reservoir configuration, can be at the origin of fault reactivation and, consequently, induced seismicity.
Two main mechanisms have been identified as responsible for grain reorganization, contingent on the initial state of stress. It is necessary to distinguish between two categories of chain forces: the unstable chain force and the stable chain force. The unstable chain forces provide support for the force thanks to the cementation. However, these structures collapse when the debonding occurs. In contrast, the stable chain forces remain in place even in the absence of cementation. It appears that the unstable chain forces are dominant in the case $k_{0}^{init}\leq k_{0}^{attr}$ and the stable chain forces are dominant in the case $k_{0}^{init}\geq k_{0}^{attr}$. In the latter case, grain reorganization occurs with grain softening.
Acknowledgements
@@
Author Contributions
All authors contributed to the study Conceptualization, Methodology, Writing (original draft), and Writing (review & editing). Software, Data curation, Investigation, Visualization: Alexandre Sac-Morane. Funding acquisition, Supervision: Hadrien Rattez, Manolis Veveakis. All authors read and approved the final manuscript.
@@ Appendix A The lognormal distribution
It appears this distribution reproduces accurently experimental observations [53]. Indeed, smaller bond surfaces and larger ones can be considered, both are crucial to the mechanical behavior of the sample. Examples of lognormal distribution used in this paper are given in Figure 23.
@@
Appendix B Bond breakage during the initialization
Table 5 shows the percentage of broken bonds (in comparison to the initial number of them) between the steps f (creation of the bonds) and h (loading at $P_{conf}$ and at $k_{0}^{init}$) of Figure 4. It is inevitable that some bonds will be broken as a consequence of the initialization algorithm, see Figure 4. Indeed, the objective is to generate a cemented granular material at a specified pressure $P_{cementation}$. Subsequently, the pressure increases to the confining value $P_{confinement}$. This increase is the source of some bond breakage (even prior to the test commencement). However, the discrepancy between the two pressures is crucial. For the sake of argument, let us assume that $P_{cementation}=P_{confinement}$. The sample is generated, and the bonds are created. The sample is stable before the bond generation process, and thus, there are no unstable chain forces. The softening of the grains exerts a dominant influence on the evolution of the sample. Moreover, the objective is to reproduce the experiments. In an oedometric test with acid injection (bond dissolution), the sample is initially cemented, set in the oedometer, loaded at confining pressure, and subsequently evolves with the injection of the acid (bond dissolution) [43]. In the experimental context, the step of applying the confining pressure also results in bond breakage.
It appears Table 5 there are more broken bonds for $k_{0}^{init}=0.2\leq k_{0}^{attr}$ than for $k_{0}^{init}=0.7\text{ or }1.5\geq k_{0}^{attr}$. During the initialization, the sample aims to reach the value $k_{0}=k_{0}^{attr}$. In the case, $k_{0}^{init}=0.2$ ($\leq k_{0}^{attr}$), the lateral dimension of the sample needs to extend. It is crucial to remember that confining pressure is still applied at the top wall. The combination of the two loadings induces mechanical bond breakage, see Equation 23.
Test characteristics N initial N final Broken (%)
$P_{conf}=1$ MPa, $k_{0}^{init}=0.2$, 2T 892 588 34
$P_{conf}=1$ MPa, $k_{0}^{init}=0.2$, 2MB 6192 5855 5
$P_{conf}=1$ MPa, $k_{0}^{init}=0.2$, 11BB 6681 6379 5
$P_{conf}=1$ MPa, $k_{0}^{init}=0.2$, 13BT 6614 6357 4
$P_{conf}=1$ MPa, $k_{0}^{init}=0.7$, 2T 899 712 21
$P_{conf}=1$ MPa, $k_{0}^{init}=0.7$, 2MB 6185 5928 4
$P_{conf}=1$ MPa, $k_{0}^{init}=0.7$, 11BB 6802 6472 5
$P_{conf}=1$ MPa, $k_{0}^{init}=0.7$, 13BT 6744 6476 4
$P_{conf}=1$ MPa, $k_{0}^{init}=1.5$, 2T 944 784 17
$P_{conf}=1$ MPa, $k_{0}^{init}=1.5$, 2MB 6104 5754 6
$P_{conf}=1$ MPa, $k_{0}^{init}=1.5$, 11BB 6709 6258 7
$P_{conf}=1$ MPa, $k_{0}^{init}=1.5$, 13MB 6664 6590 1
$P_{conf}=1$ MPa, $k_{0}^{init}=1.5$, 13BT 6817 6635 3
$P_{conf}=10$ MPa, $k_{0}^{init}=0.2$, 11BB 6973 3297 53
$P_{conf}=10$ MPa, $k_{0}^{init}=0.2$, 13MB 6867 3855 44
$P_{conf}=10$ MPa, $k_{0}^{init}=0.2$, 13BT 6988 4041 42
$P_{conf}=10$ MPa, $k_{0}^{init}=0.7$, 2T 945 423 55
$P_{conf}=10$ MPa, $k_{0}^{init}=0.7$, 2MB 6487 3659 44
$P_{conf}=10$ MPa, $k_{0}^{init}=0.7$, 11BB 6847 3977 42
$P_{conf}=10$ MPa, $k_{0}^{init}=0.7$, 13MB 6986 5218 25
$P_{conf}=10$ MPa, $k_{0}^{init}=0.7$, 13BT 7079 4889 31
$P_{conf}=10$ MPa, $k_{0}^{init}=1.5$, 2T 913 392 57
$P_{conf}=10$ MPa, $k_{0}^{init}=1.5$, 2MB 6438 3205 50
$P_{conf}=10$ MPa, $k_{0}^{init}=1.5$, 11BB 6886 3924 43
$P_{conf}=10$ MPa, $k_{0}^{init}=1.5$, 13MB 6888 4565 34
$P_{conf}=10$ MPa, $k_{0}^{init}=1.5$, 13BT 7084 4347 39
Table 5: Percentage of bonds broken during the initialization (steps f to h of Figure 4). Only the tests with a percentage $\geq 1\%$ are listed.